# Neural refractive index field: Unlocking the Potential of Background-oriented Schlieren Tomography in Volumetric Flow Visualization

### v1

@@ 4.1 Numerical validation
4.2 Experimental validation
The numerical validation presented above has confirmed the robustness of the NeRIF technique. Now, we provide an experimental demonstration of this approach. Our study focused on a turbulent premixed methane/air flame with an equivalence ratio of $\phi=1.2$. Within the reactor’s central nozzle, depicted in Figure 2, the fuel-air mixture was ignited amidst a nitrogen co-flow. Utilizing the NeRIF technique, we reconstructed the temporal evolution of the three-dimensional density field, as illustrated in Figure 5(a). This reconstruction effectively circumvented the overlap of flame surface structures, enabling simultaneous visualization of the flame front leading edge and the interface between the hot combustion products and the nitrogen co-flow.
Given the absence of ground truth data for the direct comparison, we validated our reconstruction with an alternative approach. Out of the nine projections captured, eight were employed for the reconstruction process and one was saved for the validation. Figures 5(b-c) showcase a comparison between the 9-th measured projection and the re-projection simulated with the reconstructed refractive index field, in terms of metrics such as SSIM, PSNR, and CC.
For eight angles employed in the reconstruction process, the PSNR consistently surpasses 30 when comparing the measured projections with their corresponding re-projections across different phantoms. Furthermore, the SSIM exceeds 0.95, and the CC stands above 0.99, demonstrating a remarkable level of fidelity in the reconstructed three-dimensional density field. We also provide a video demonstration of the reconstructed 3D density field from 0 to 400 ms, available in Movie 1. As a result, the reconstruction can be considered as a highly reliable algorithm and utilized for further investigations on the intricate mechanism of turbulent flames.
5 Conclusions
@@
In summary, the NeRIF signifies a noteworthy advancement in BOST, offering great prospects for three-dimensional flow and combustion diagnostics applications. More importantly, though demonstrated in the context of BOST, the core mechanisms of NeRIF are readily adaptable to diverse tomographic modalities, including tomographic absorption spectroscopy and tomographic particle imaging velocimetry, thereby extending its potential reach across various domains of the fluid visualization and analysis.
[…]
@@
[Funding]This work was funded by National Science Foundation of China under grant number 51706141, 51976122 and SJTU Global Strategic Partnership Fund (2023 SJTU-UoE).
[…]
[Declaration of interests]The authors declare that there is no conflict of interests regarding the publication of this article.
[…]
[Author ORCIDs]YH, https://orcid.org/0009-0007-7393-1032.
[…]
[Acknowledgement]We would like acknowledge Prof. Samuel J. Grauer for the valuable suggestions.

### v2

@@ IV.2 Experimental validation
Figure 8: Visualization of X-axis and Y-axis slices of flame at 0 ms
V Conclusions
@@
In summary, the NeRIF signifies a noteworthy advancement in BOST, offering great prospects for three-dimensional flow and combustion diagnostics applications. More importantly, though demonstrated in the context of BOST, the core mechanisms of NeRIF are readily adaptable to diverse tomographic modalities, including tomographic absorption spectroscopy and tomographic particle imaging velocimetry, thereby extending its potential reach across various domains of fluid visualization and analysis.
Acknowledgements.
@@
This work was funded by National Science Foundation of China under grant number 51706141, 51976122 and SJTU Global Strategic Partnership Fund (2023 SJTU-UoE). We would like acknowledge Prof. Samuel J. Grauer for the valuable suggestions.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Structure of Neural Network
The structure of the neural network in this study can be shown in the following table.
Table 6: Structure of the Neural Network.
Layers Dimension
Input layer $3\times(2N+1)$
Layer 1 256
Layer 2 256
Layer 3 256
Layer 4 256+$3\times(2N+1)$
Layer 5 256
Layer 6 256
Layer 7 256
Layer 8 64+64+64+64
Output layer 1+1+1+1
The four outputs are the refractive index $n$ and its gradient $\nabla n$ in directions in x, y and z. Here we compare the performance of directly fully connected neural network and our adopted network.In this study, the influence of the last layer structure is tested experimentally, and the convergence characteristics of the Multi-head NN used in this paper are compared with the multi-layer perceptron MLP connected by jumps. As shown in Fig. 9, the structure we adopted has better convergence characteristics, with the mean value of loss being 0.07 and the mean value of MLP loss being 0.8.
Figure 9: The influence of neural network structure on reconstruction
The positional encoding formula presented in Eq. 8 is designed to inject relative or absolute position information into the model, where N is the parameter selected according to the complexity of the reconstructed field. Such a coding method can avoid excessive smoothing caused by implicit regularization of the neural network so that the neural network can fit the high-frequency information of the flow field in the early stage, and improve the convergence speed and accuracy. As shown in Fig. 10, it can be seen that due to the function of coordinate coding, loss decreases slightly slowly in the initial stage of reconstruction, but the speed of decrease is more uniform and the expected loss threshold can be reached faster in the end, but it is worth noting that excessive coordinate coding may also lead to overfitting of the refractive index field.
$E(p)=\left(\sin\left(2^{0}\pi{p}\right),\cos\left(2^{0}\pi{p}\right),\dots,\sin\left(2^{N-1}\pi{p}\right),\cos\left(2^{N-1}\pi{p}\right)\right)$ (8)
Figure 10: Effect of position encoding on reconstruction
Appendix B Ray Tracing for Simulation
During the simulation, we used the fourth-order accuracy of the Runge-Kutta method for ray tracing from the camera plane to obtain the propagation of the rays in a non-uniform refractive index field and the displacement of the camera rays.We introduce the position matrix $R$ and the ray vector $T$:
$R=\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ (9)
$T=\frac{dr}{dt}=n(r)\frac{dr}{ds}=n(r)\begin{pmatrix}\cos\alpha\\ \cos\beta\\ \cos\gamma\end{pmatrix}$ (10)
Define the matrix $D$:
$D=n(r)\nabla n(r)=\frac{1}{2}\nabla n(r)^{2}=\frac{1}{2}\begin{pmatrix}\frac{\partial n^{2}}{\partial x}\\ \frac{\partial n^{2}}{\partial y}\\ \frac{\partial n^{2}}{\partial z}\end{pmatrix}$ (11)
Combining equations, we get:
$T^{\prime}=D(r)$ (12)
Using the Runge-Kutta method, we compute:
$K_{1}$ $=\Delta tD(R_{n})$ (13)
$K_{2}$ $=\Delta tD\left(R_{n}+\frac{\Delta t}{2}T_{n}+\frac{\Delta t}{8}K_{1}\right)$
$K_{3}$ $=\Delta tD\left(R_{n}+\frac{\Delta t}{2}T_{n}+\frac{\Delta t}{2}K_{2}\right)$
$K_{4}$ $=\Delta tD\left(R_{n}+\frac{\Delta t}{2}T_{n}+\frac{\Delta t}{2}K_{3}\right)$
$R_{n+1}$ $=R_{n}+\frac{\Delta t}{6}(6T_{n}+K_{1}+2K_{2}+2K_{3}+K_{4})$
$T_{n+1}$ $=T_{n}+\frac{\Delta t}{6}(K_{1}+4K_{2}+K_{3}+K_{4})$
Here, $\Delta t$ is the step size. The Runge-Kutta method computes the coordinates and direction of the ray iteratively, obtaining multiple coordinate points along the ray’s path. This method does not require consideration of the direction and position at each refractive index gradient boundary. Connecting these points easily reveals the ray’s trajectory through the gradient index medium. As an example, the displacement obtained by tracing phantom2 is shown in Fig. 11.
Figure 11: Projected displacement of spray combustion obtained by simulation
Appendix C Image Quality Assessment Metrics
We use three image quality assessment metrics in this study: Structural Similarity Index Measure (SSIM), Correlation Coefficient (CC), and Peak Signal-to-Noise Ratio (PSNR). These metrics are widely used in the field of image processing to evaluate the quality of reconstructed images.
C.1 Structural Similarity Index Measure (SSIM)
The Structural Similarity Index Measure (SSIM) is used for measuring the similarity between two images. It is a perception-based model that considers image degradation as a perceived change in structural information, while incorporating important perceptual phenomena, including both luminance masking and contrast masking terms.
$\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+c_{1})(2\sigma_{xy}+c_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2})}$ (14)
In Eq. 14, $\mu_{x}$ and $\mu_{y}$ are the averages of $x$ and $y$ respectively, $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are the variances, and $\sigma_{xy}$ is the covariance. The constants $c_{1}$ and $c_{2}$ are used to stabilize the division with weak denominator.
C.2 Correlation Coefficient (CC)
The Correlation Coefficient is a statistical measure that calculates the strength of the linear relationship between two variables. It is dimensionless and varies between -1 and +1, where 1 means a perfect positive linear relationship, -1 means a perfect negative linear relationship, and 0 means no linear relationship at all.
$\text{CC}(X,Y)=\frac{\sum(X_{i}-\bar{X})(Y_{i}-\bar{Y})}{\sqrt{\sum(X_{i}-\bar{X})^{2}\sum(Y_{i}-\bar{Y})^{2}}}$ (15)
C.3 Peak Signal-to-Noise Ratio (PSNR)
PSNR is most commonly used to measure the quality of reconstruction of lossy compression codecs (e.g., for image compression). The signal in this case is the original data, and the noise is the error introduced by compression.
$\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}_{I}^{2}}{\text{MSE}}\right)$ (16)
Here, $\text{MAX}_{I}$ is the maximum possible pixel value of the image, and MSE stands for Mean Squared Error between the original and compressed image.
Appendix D Spatial Resolution Determination
There are various ways to define spatial resolution (SR). In this work, we adopt a definition based on the modulation transfer function (MTF) (Yu et al., 2017). The line spread function (LSF) is calculated by taking the first derivative of the edge spread function (ESF). The MTF is then obtained by performing a 1D discrete Fourier transform of the LSF. In the numerical experiments, we constructed a piecewise function within a 4 $\times$ 4 $\times$ 4 unit space, as shown in Eq. 17. The derivative of this function features sharp boundaries, where the density gradient at the boundary surface is a step function, which can be used for MTF calculation. A 10% cutoff threshold was applied to the MTF, resulting in an SR of 61.58 lp/unit,The distribution of 3D spatial resolution is shown in Fig. 12.
$n(x,y,z)=\begin{cases}1.5-\left(|x|+|y|+|z|\right),&\text{if }|x|+|y|+|z|\leq 1.5\\ 0,&\text{otherwise}\end{cases}$ (17)
Figure 12: Distribution of three-dimensional spatial resolution